UT 20130522 Technical Observing - MODS Collimation Issues

Observer: DMiller, JHill, OKuhn (Tucson) ASkemer, VBailey (LBTO)
Telescope Operator: GBechetti (LBTO)
AO Support: JCGuerra (Home)
Telescope Support: (LBTO)
Instrument Support: (LBTO)

Summary

LFBG

03:00 Primary temperature gradients. See Z11, Z22 etc. WFSC looks like donut. Missing two left

03:30 Turn on high order corrections. Too much aberration so converged to poor image. Clear Primary Act Opt. Helps a bit. Clear primary and secondary. Collimates well

04:18 Turn off High order corrections. Leave current high order correction on to see stability. Stable. Getting down to 48 EL. Move higher

04:32 preset, BS9152, GS=1, TS=1. Primary mirror panic. Outer loop stalled message from PMC.

04:45 preset. guider image is awful! Active Optics corrects. Collimates

04:55 mods1b.20130522.00005/6.fits 10/30 second exposures. 10s image has structure. 30s has fairly round images.

05:00 preset, BS9152, GS=1, TS=5. image 7/8 30/10 second still elongation x = 6.7, y = 9.8 pixel fwhm.

05:20 setFocus left 5. (from 0.0) get 611 + 192 + 164 nm. MODS image #9 Seeing a worse, but more round images

05:30 setFocus left -5. (from 0.0) MODS 10 still elongated

05:50 preset BS9152, GS=8, TS=2, /YAXIS Trefoil converged value change by 800nm, change to 9-2 recon. Stable

06:35 preset BS9152, GS=8, TS=2, PA=160 get x and y axis. Stable.

More tests that show the new reconstructors do not work, even with no vignetting. Back to 20130118 reconstructor.

-- DougMiller - 22 May 2013
Topic revision: r1 - 22 May 2013, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback